OpenStudy (pphalke):

I need help I feel like no one is helping me and I looked it up too The balance on Jim's credit card on April 17, the billing date, was $78.30. For the period ending May 16, he made the following transactions. April 20 payment $50.00 May 1 charge for gas: $29.20 a. find the average daily balance for the billing period b. find the finance charge that is due may 17 c. find the balance that is due may 17

OpenStudy (legomyego180):

Not sure.

OpenStudy (pphalke):

anyone good at this?

OpenStudy (pphalke):

no idea :(

jimthompson5910 (jim_thompson5910):

Apr 17: Balance = $78.30 Apr 18: Balance = $78.30 Apr 19: Balance = $78.30 Apr 20: Balance = $28.30 ... payment made (78.30-50=28.30) Apr 21: Balance = $28.30 Apr 22: Balance = $28.30 Apr 23: Balance = $28.30 Apr 24: Balance = $28.30 Apr 25: Balance = $28.30 Apr 26: Balance = $28.30 Apr 27: Balance = $28.30 Apr 28: Balance = $28.30 Apr 29: Balance = $28.30 Apr 30: Balance = $28.30 May 1: Balance = $57.50 ... charge (28.30+29.20 = 57.5) May 2: Balance = $57.50 May 3: Balance = $57.50 May 4: Balance = $57.50 May 5: Balance = $57.50 May 6: Balance = $57.50 May 7: Balance = $57.50 May 8: Balance = $57.50 May 9: Balance = $57.50 May 10: Balance = $57.50 May 11: Balance = $57.50 May 12: Balance = $57.50 May 13: Balance = $57.50 May 14: Balance = $57.50 May 15: Balance = $57.50 May 16: Balance = $57.50 There are 3 days (apr 17 through apr 19) where the balance is 78.30 There are 11 days (apr 20 though apr 30) where the balance is 28.30 There are 16 days (may 1 through may 16) where the balance is $57.50 Let me know if this helps with calculating the average daily balance

OpenStudy (pphalke):

it helps that way what is the daily average balance for the billing period? how do I find that

OpenStudy (skullpatrol):

What does the word "average" mean to you?

OpenStudy (pphalke):

adding

OpenStudy (skullpatrol):

Nothing but adding?

jimthompson5910 (jim_thompson5910):

multiply the number of days with the corresponding balance for apr 17 through apr 20 3*78.30 = 234.9 for apr 20 through apr 30 11*28.30 = 311.3 for may 1 through may 16 16*57.50 = 920 Now you'll add up the products (234.9 and 311.3 and 920) Divide that sum by the total number of days from apr 17 to may 16

jimthompson5910 (jim_thompson5910):

you could take the slower way and actually add all of the 30 values up and divide by 30, but the weighted mean approach is much faster

OpenStudy (skullpatrol):

Recall that the "average" of a set of numbers is the sum of the numbers divided by the number of numbers.

OpenStudy (pphalke):

$48.87

jimthompson5910 (jim_thompson5910):

correct, @pphalke

jimthompson5910 (jim_thompson5910):

that's the average daily balance for that 30 day billing cycle

OpenStudy (pphalke):

then I got $0.21 for the next answer I don't know how I did that

jimthompson5910 (jim_thompson5910):

to find the finance charge (interest), use this formula i = P*r*t i = interest = finance charge P = balance (in this case, average daily balance) r = interest rate t = time in years (1 month = 1/12 years)

jimthompson5910 (jim_thompson5910):

i = P*r*t i = 48.87*r*(1/12) I don't know what the interest rate is so I can't continue